OpenStudy (anonymous):

Statistics Probability: 4 people are at a car showroom. They may purchase the car (denoted as P) or may not (denoted as N). What is the probability that (A) exactly 3 people will purchase a car, (B) 3 or fewer people will purchase a car, (C) 1 or more people will purchase a car, and (D) all four people make the same purchase decision?

OpenStudy (anonymous):

What is the probability that they purchase the car? Is it equally probable that they do or do not buy it?

OpenStudy (anonymous):

they are equally liekly to purchase the vehicle

OpenStudy (anonymous):

Okay. Now, the events that someone buys the car is independent of whether or not other people buy the car. So for each person who buys a car, you multiply \(\large \frac{1}{2}\).

OpenStudy (anonymous):

Also, the event that 2 people buy a car and the event that more or less than 2 people buy a car can not happen at the same time. They are mutual exclusive events. So to find the probability that either happen, you add them up.

OpenStudy (anonymous):

So for example, the probability that less than 3 people buy the car is the probability that 1 person buys a car OR that 2 people buy a car. So we have \(\frac{1}{2} + \frac{1}{2} \cdot \frac{1}{2}\)

OpenStudy (anonymous):

Are you following me so far?

OpenStudy (anonymous):

One last thing to remember is that probability that an event DOESN'T happen is just one minus the probability that it does. So the probability that you don't have 3 people buy the vehicle is \(1 - \frac{1}{2} \cdot \frac{1}{2} \cdot \frac{1}{2}\).